# Convergence and long-time behavior of finite volumes for a generalized Poisson-Nernst-Planck system with cross-diffusion and size exclusion

### v1

@@ Abstract.
We present a finite volume scheme for modeling the diffusion of charged particles, specifically ions, in constrained geometries using a degenerate Poisson-Nernst-Planck system with size exclusion yielding cross-diffusion. Our method utilizes a two-point flux approximation and is part of the exponentially fitted scheme framework. The scheme is shown to be thermodynamically consistent, as it ensures the decay of some discrete version of the free energy. Classical numerical analysis results – existence of discrete solution, convergence of the scheme as the grid size and the time step go to 00 – follow. We also investigate the long-time behavior of the scheme, both from a theoretical and numerical point of view. Numerical simulations confirm our findings, but also point out some possibly very slow convergence towards equilibrium of the system under consideration.
@@ 1. Presentation of the problem
1.1. The continuous generalized Poisson-Nernst-Planck model
Motivated by the transfer of ions in confined geometries, Burger et al. introduced in [5] a model accounting for cross-diffusion and size-exclusion effects. The model [6] further incorporated self-consistent electric interaction. In this model, $I$ species, the volume fractions of which being denoted by $U=(u_{i})_{1\leq i\leq I}$, are subject to diffusion as well as to electric forces induced by a self-consistent electrostatic potential. Denote by $\Omega\subset\mathbb{R}^{d}$ a bounded connected polyhedral domain, then the conservation of the volume occupied by the species $i$ writes
@@ 1.1. The continuous generalized Poisson-Nernst-Planck model
shall be thought of as the volume fraction of available space for the ions, possibly occupied by a motile and electro-neutral solvent. The quantity $u_{0}$ is then required to remain nonnegative, leading to size exclusion for the other species $u_{i}$, $i=1,\dots,I$. Denoting by $z_{i}\in\mathbb{Z}$ the charge of species $i$ and by $\lambda>0$ the (scaled) Debye length, then the electrostatic potential solves the Poisson equation
(4) $-\lambda^{2}\Delta\phi=\sum_{i=1}^{I}z_{i}u_{i}+f$
for some prescribed background charge density $f$. We consider boundary conditions of mixed type for the electric potential. More precisely, we assume that the boundary $\partial\Omega$ of the domain can be split into a insulator part $\Gamma^{N}$ and its complement $\Gamma^{D}$ on which Dirichlet boundary condition is imposed:
(5) $\nabla\phi\cdot n=0\quad\text{on}\;\Gamma^{N}\qquad\text{and}\quad\phi=\phi^{D}\;\text{on}\;\Gamma^{D}.$
Throughout this paper, we will assume that $f\in L^{\infty}(\Omega)$ and that $\phi^{D}$ is the trace of an $L^{\infty}\cap H^{1}(\Omega)$ function (which we also denote by $\phi^{D}$). Neither $f$ nor $\phi^{D}$ depend on time. Boundary conditions of various types can be considered for the conservation laws (1)–(2), like for instance Robin type boundary condition modeling electrochemical reaction thanks to Butler-Volmer type formula, see for instance [8], or boundary conditions of mixed Dirichlet-Neumann type as in [28]. In the presentation of the scheme, we assume for simplicity that the system is isolated, in the sense that
@@
$\mathcal{A}=\left\{U=\left(u_{i}\right)_{1\leq i\leq I}\in\mathbb{R}_{+}^{I}\;\middle|\;\sum_{i=1}^{I}u_{i}\leq 1\right\}$
@@ 1.2. Entropy structure of the model
Let us now described the entropy (or formal gradient flow) structure of the model. Introduce the Slotboom variables $w_{i}=\frac{u_{i}}{u_{0}}e^{z_{i}\phi}$, then the fluxes (2) rewrite as
@@
Multiplying (1) by the so-called electrochemical potentials $\mu_{i}=\log w_{i}=\log\frac{u_{i}}{u_{0}}+z_{i}\phi$, integrating over $\Omega$ and summing over $i=1,\dots,I$ yields
@@
Assume that the $u_{i}$ are positive for $i=0,\dots,I$ (as proved in the discrete case later on), then the second term in (9) is well-defined and non-negative. As a consequence, the free energy decays along time, as a manifestation of the second principle of thermodynamics. Observe that $\mathcal{H}$ need not be non-negative but may be bounded uniformly from below by a constant depending only on $\lambda$, $f$ and $\phi^{D}$.
@@ 1.3. Weak solution
(12) $0\leq u_{i}\leq 1$ for $i=0,\dots,I$ and a.e. $(t,x)\in\mathbb{R}_{+}\times\Omega$.
@@
(13) $\|\phi\|_{L^{\infty}(\mathbb{R}_{+}\times\Omega)}+\|\phi-\phi^{D}\|_{L^{\infty}(\mathbb{R}_{+};V)}\leq C(\lambda,z,\Omega,f),$
@@
is equipped with the $H^{1}(\Omega)$ semi-norm $\|\nabla v\|_{L^{2}}$, which is a norm thanks to Poincaré inequality. We have emphasized in the right-hand side of (13) the dependence of the bound on the data, especially on the Debye length $\lambda$. The estimate (9) is the other main estimate on which our study will rely. It provides enough control to define a proper notion of weak solutions, but at the price of a suitable reformulation of the fluxes (2). Let us first remark that, since the free energy $\mathcal{H}$ is bounded (see Lemma 3.6), integrating (9) over $t\in\mathbb{R}_{+}$ provides
(14) +×$\iint_{\mathbb{R}_{+}\times\Omega}\sum_{i=1}^{I}D_{i}u_{0}^{2}e^{-z_{i}\phi}|\nabla\sqrt{w_{i}}|^{2}\leq C.$
As already noticed in [28], the above inequality yields a control in $L^{2}_{\text{loc}}(\mathbb{R}_{+};H^{1}(\Omega))$ on $\sqrt{u_{0}}$ and on $u_{0}$ itself, as well as some control on some product terms involving $u_{0}$ and $u_{i}$ for $i\geq 1$. From the numerical analysis exposed hereafter, we derive a $L^{2}_{\text{loc}}(\mathbb{R}_{+};H^{1}(\Omega))$ estimate on $u_{i}\sqrt{u_{0}}$, which, together with (12) provides a $L^{2}_{\text{loc}}(\mathbb{R}_{+};H^{1}(\Omega))$ control on $u_{i}u_{0}$ too. Hence, all the terms in the following expression of the fluxes
@@
  • $u_{i}\in L^{\infty}(\mathbb{R}_{+}\times\Omega)$ with $U\in\mathcal{A}$ a.e. in +×$\mathbb{R}_{+}\times\Omega$, with moreover $\sqrt{u_{0}}$ and $u_{i}\sqrt{u_{0}}$ belonging to $L^{2}_{\text{loc}}(\mathbb{R}_{+};H^{1}(\Omega))$ for $u_{0}$ defined as in (3);
  • $\phi\in L^{\infty}(\mathbb{R}_{+}\times\Omega)$ with $\phi-\phi^{D}\in L^{\infty}(\mathbb{R}_{+};V)$ satisfies
@@
  • for all $\varphi\in C^{1}_{c}(\mathbb{R}_{+}\times\overline{\Omega})$ and all $i=1,\dots,I$, there holds
    (17) +×$\iint_{\mathbb{R}_{+}\times\Omega}u_{i}\partial_{t}\varphi+\int_{\Omega}u_{i}^{0}\varphi(0,\cdot)-\iint_{\mathbb{R}_{+}\times\Omega}D_{i}\big{(}\nabla(u_{0}u_{i})-4u_{i}\sqrt{u_{0}}\nabla\sqrt{u_{0}}+u_{i}u_{0}z_{i}\nabla\phi\big{)}\cdot\nabla\varphi=0.$
1.4. Equilibrium and modified Poisson-Boltzmann equation
@@
where $\boldsymbol{\xi}=\left(\xi_{j}\right)_{1\leq j\leq I}\in\mathbb{R}^{I}$. Then the densities $U^{\infty}=\left(u_{i}^{\infty}\right)_{1\leq i\leq I}$ satisfy
@@
Plugging these expressions in the Poisson equation (4) provides what we call the modified Poisson-Boltzmann equation, i.e.
(20) $-\lambda^{2}\Delta\phi^{\infty}+r(\phi^{\infty},\boldsymbol{\mu}^{\infty})=f\quad\text{in}\;\Omega,$
complemented with the same boundary conditions (5). In (20), $r$ is the smooth function from ×$\mathbb{R}\times\mathbb{R}^{I}$ to $\mathbb{R}$ defined by
@@
$\partial_{y}r(y,\boldsymbol{\xi})=\sum_{i=0}^{I}(z_{i})^{2}v_{i}(y,\boldsymbol{\xi})-\left(\sum_{i=0}^{I}z_{i}v_{i}(y,\boldsymbol{\xi})\right)^{2}\geq 0,$
the last inequality being a consequence of the convexity of $y\mapsto y^{2}$ as $\sum_{i=0}^{I}v_{i}(y,\boldsymbol{\xi})=1$. Therefore $r$ is non-decreasing w.r.t. its first variable $y$, ensuring the well-posedness of the problem (20)&(5). Getting one step further allows to check that the solution $(\phi^{\infty},\boldsymbol{\mu}^{\infty})$ is the unique minimizer of the convex functional $\Psi:(\phi^{D}+V)\times\mathbb{R}^{I}\to\mathbb{R}$ defined by
@@
Despite hints in this directions, the uniqueness of the steady state is not established in [6], and no quantitative convergence result has been established so far.
@@ 1.5. Goal and positioning of the paper
The goal of this paper is to provide a developed mathematical study of a numerical scheme that was proposed in our previous contribution [11]. It enters the broad family of the two-point flux approximation (TPFA) finite volume schemes [23], which has been applied to several cross-diffusion problems in the recent years (see e.g. [10, 15, 33, 27, 13, 34, 9]). Our scheme mixes ideas from the so-called square-root approximation (SQRA) finite volume scheme [12] (see also [35, 30]) with some features of the Scharfetter-Gummel scheme [36, 18], especially the use of the Bernoulli function (28) in the definition of the fluxes. Note that a purely SQRA approach was also proposed in [11], but since it is overpassed by the mixed approach both in terms of accuracy and of robustness in the small Debye length (or quasi-neutral) limit, we do not develop here the analysis. Let us however stress that our whole study extends to this full-SQRA scheme at the price of very minor modifications sketched in [11].
@@
As highlighted in [11], our scheme fulfils key properties. First, it is second order accurate in space, in opposition to the scheme proposed in [7] in which upwinding lead to a mere first order consistency. Even though our scheme does not use the so-called entropy variables (the electrochemical potentials in our framework) as unknowns in opposition to the finite element scheme presented in [29], our scheme satisfies a discrete energy dissipation inequality, which should be thought as a discrete counterpart to (9), and which is key for its numerical analysis. Keeping conservative variables as primary variable yields well-behaved nonlinear solvers (we used Newton-Raphson), allowing to consider larger time steps than for the finite element methods of [29] or for the finite volume method of [2]. Together with the stability to be used for proving the convergence of the scheme, the discrete energy dissipation inequality implies that the scheme preserves exactly the steady states of the continuous model in the sense that, at the discrete level, fluxes vanish when concentrations are of the form (19). The unique steady state of our scheme is then characterized only by the electric potential which solves a discrete counterpart of the Poisson–Boltzmann equation (20), which can be computed directly in an efficient way.
@@
In the present contribution, we first rigorously establish the convergence of the scheme when the discretization parameters (mesh size and time step) tend to 00. As in the continuous setting [28], the proof is based on compactness arguments. It strongly relying on the discrete counterpart of the energy dissipation estimate (9), following the boundedness-by-entropy method [32], and more precisely a discrete version of it [16, 34]. Uniqueness of the weak solutions is moreover established in [28] under moderate regularity assumptions thanks to the uniqueness technique of Gajewski [24, 25].
@@
We also investigate the long time behaviour of the numerical scheme. We show in particular that the discrete solution to our scheme for the evolution problem converges as time goes to infinity towards the unique solution to the discrete steady problem. As the proof relies on LaSalle’s invariance principle, no quantitative convergence rate is derived. Numerical simulations suggest that, the convergence rate strongly depends on the Debye length, but also on the initial profile $U^{0}$ (in opposition to the classical Poisson-Nernst-Planck problem without volume filling [3]).

### v2

# Convergence and long-time behavior of finite volumes for a generalized Poisson–Nernst–Planck system with cross-diffusion and size exclusion

@@ Abstract.
We present a finite volume scheme for modeling the diffusion of charged particles, specifically ions, in constrained geometries using a degenerate Poisson–Nernst–Planck system with size exclusion yielding cross-diffusion. Our method utilizes a two-point flux approximation and is part of the exponentially fitted scheme framework. The scheme is shown to be thermodynamically consistent, as it ensures the decay of some discrete version of the free energy. Classical numerical analysis results – existence of a discrete solution, convergence of the scheme as the grid size and the time step go to 0 – follow. We also investigate the long-time behavior of the scheme, both from a theoretical and numerical point of view. Numerical simulations confirm our findings, but also point out some possibly very slow convergence towards equilibrium of the system under consideration.
@@ 1. Presentation of the problem
1.1. The continuous generalized Poisson–Nernst–Planck model
Motivated by the transfer of ions in confined geometries, Burger et al. introduced in [5] a model accounting for cross-diffusion and size-exclusion effects. The model [6] further incorporated self-consistent electric interaction. In this model, $I$ species, the volume fractions of which being denoted by $U=(u_{i})_{1\leq i\leq I}$, are subject to diffusion as well as to electric forces induced by a self-consistent electrostatic potential. Denote by $\Omega\subset\mathbb{R}^{d}$ a bounded connected polyhedral domain, then the conservation of the volume occupied by the species $i$ is written
@@ 1.1. The continuous generalized Poisson–Nernst–Planck model
shall be thought of as the volume fraction of available space for the ions, possibly occupied by a motile and electro-neutral solvent. The quantity $u_{0}$ is then required to remain non-negative, leading to size exclusion for the other species $u_{i}$, $i=1,\dots,I$. Denoting by $z_{i}\in\mathbb{Z}$ the charge of species $i$ and by $\lambda>0$ the (scaled) Debye length, then the electrostatic potential $\phi$ solves the Poisson equation
(4) $\lambda^{2}\Delta\phi=\sum_{i=1}^{I}z_{i}u_{i}+f$
for some prescribed background charge density $f$. We consider boundary conditions of mixed type for the electric potential. More precisely, we assume that the boundary $\partial\Omega$ of the domain can be split into an insulator part $\Gamma^{N}$, which is a relatively open $C^{1}$ part of the boundary, and its complement $\Gamma^{D}$ on which a Dirichlet boundary condition is imposed:
(5) $\nabla\phi\cdot n=0\quad\text{on}\;\Gamma^{N}\qquad\text{and}\quad\phi=\phi^{D}\;\text{on}\;\Gamma^{D},$
where $n$ is the unit normal to $\partial\Omega$. Throughout this paper, we will assume that $f\in L^{\infty}(\Omega)$ and that $\phi^{D}$ is the trace of an $L^{\infty}\cap H^{1}(\Omega)$ function (which we also denote by $\phi^{D}$). Neither $f$ nor $\phi^{D}$ depend on time. Boundary conditions of various types can be considered for the conservation laws (1)–(2), for instance a Robin type boundary condition modeling electrochemical reaction thanks to a Butler–Volmer type formula, see for instance [8], or boundary conditions of mixed Dirichlet–Neumann type as in [31]. In the presentation of the scheme, we assume for simplicity that the system is isolated, in the sense that
@@
$\mathcal{A}=\left\{U=\left(u_{i}\right)_{1\leq i\leq I}\in({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}})^{I}\;\middle|\;\sum_{i=1}^{I}u_{i}\leq 1\right\}$
@@ 1.2. Entropy structure of the model
Let us now describe the entropy (or formal gradient flow) structure of the model. Introduce the Slotboom variables $w_{i}=\frac{u_{i}}{u_{0}}e^{z_{i}\phi}$, then the fluxes (2) rewrite as
@@
If the so-called electrochemical potentials defined by $\mu_{i}=\log w_{i}=\log\frac{u_{i}}{u_{0}}+z_{i}\phi$ are regular enough, say in $H^{1}\cap L^{\infty}(\Omega)$, then $\nabla\log w_{i}\cdot\nabla w_{i}=4|\nabla\sqrt{w_{i}}|^{2}.$ Therefore, multiplying (1) by $\mu_{i}$, integrating over $\Omega$ and summing over $i=1,\dots,I$ then yields
@@
Assume that the $u_{i}$ are positive for $i=0,\dots,I$ (as proved in the discrete case later on), then the second term in (9) is well-defined and non-negative. As a consequence, the free energy decays along-time, as a manifestation of the second principle of thermodynamics. Observe that $\mathcal{H}$ need not be non-negative but may be bounded uniformly from below by a constant depending only on $\lambda$, $f$, and $\phi^{D}$.
@@ 1.3. Weak solution
(12) $0\leq u_{i}\leq 1$ for $i=0,\dots,I$ and a.e. $(t,x)\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega$.
@@
(13) $\|\phi\|_{L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega)}+\|\phi-\phi^{D}\|_{L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};V)}\leq C(\lambda,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(z_{i})_{i}},\Omega,f),$
@@
[…]
(14) $\iint_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega}\sum_{i=1}^{I}D_{i}u_{0}^{2}e^{-z_{i}\phi}|\nabla\sqrt{w_{i}}|^{2}\leq C.$
As already noticed in [31], the above inequality yields a control in $L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};H^{1}(\Omega))$ on $\sqrt{u_{0}}$ and on $u_{0}$ itself, as well as control on some product terms. More precisely, in the numerical analysis exposed hereafter, we derive a $L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};H^{1}(\Omega))$ estimate on $u_{i}\sqrt{u_{0}}$, which, together with (12) provides a $L^{2}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};H^{1}(\Omega))$ control on $u_{i}u_{0}$ too. Hence, all the terms in the following expression of the fluxes
@@
[…]
  • $\phi\in L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega)$ with $\phi-\phi^{D}\in L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};V)$ satisfies
@@
  • for all $\varphi\in C^{1}_{c}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\overline{\Omega})$ and all $i=1,\dots,I$, there holds
    (17) $\iint_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega}u_{i}\partial_{t}\varphi+\int_{\Omega}u_{i}^{0}\varphi(0,\cdot)-\iint_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega}D_{i}\big{(}\nabla(u_{0}u_{i})-4u_{i}\sqrt{u_{0}}\nabla\sqrt{u_{0}}+u_{i}u_{0}z_{i}\nabla\phi\big{)}\cdot\nabla\varphi=0.$
1.4. Equilibrium and modified Poisson–Boltzmann equation
@@
where $\boldsymbol{\xi}=\left(\xi_{j}\right)_{1\leq j\leq I}\in\mathbb{R}^{I}$. Then the volume fractions $U^{\infty}=\left(u_{i}^{\infty}\right)_{1\leq i\leq I}$ satisfy
@@
Plugging these expressions in the Poisson equation (4) provides what we call the modified Poisson–Boltzmann equation, i.e.
(20) $\lambda^{2}\Delta\phi^{\infty}+r(\phi^{\infty},\boldsymbol{\mu}^{\infty})=f\quad\text{in}\;\Omega,$
complemented with the same boundary conditions (5). In (20), $r$ is the smooth function from $\mathbb{R}\times\mathbb{R}^{I}$ to $\mathbb{R}$ defined by
@@
$\partial_{y}r(y,\boldsymbol{\xi})=\sum_{i=0}^{I}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}z_{i}^{2}}v_{i}(y,\boldsymbol{\xi})-\left(\sum_{i=0}^{I}z_{i}v_{i}(y,\boldsymbol{\xi})\right)^{2}\geq 0,$
the last inequality being a consequence of the convexity of $y\mapsto y^{2}$ as $\sum_{i=0}^{I}v_{i}(y,\boldsymbol{\xi})=1$. Therefore, $r$ is non-decreasing w.r.t. its first variable $y$. The monotonicity and boundedness of $r$ allow us to conclude, via classical monotonicity arguments, the well-posedness of the problem (20)&(5). Going one step further allows us to check that the solution $(\phi^{\infty},\boldsymbol{\mu}^{\infty})$ is a minimizer of the functional $\Psi:(\phi^{D}+V)\times\mathbb{R}^{I}\to\mathbb{R}$ defined by
@@
One readily checks that $\Psi$ is strictly convex, and thus the minimiser $(\phi^{\infty},\boldsymbol{\mu}^{\infty})$ is unique. Moreover, one easily verifies that minimizers for $\Psi$ are steady states of the system (1)–(7). Despite hints in these directions, the uniqueness of the steady state is neither established in [6], nor, up to our knowledge, in more recent contributions. Furthermore, no quantitative convergence result has been established so far. Non-quantitative results were however recently presented in [19] in the absence of electrical interactions, i.e. $z_{i}=0$ for all $i$. This gap is partially filled in the discrete case. We rigorously establish in Section 5 the uniqueness of the steady states, which are minimisers of some discrete counterpart of the functional $\Psi$, cf (35), and the (non-quantitative) convergence towards such a steady profile as time goes to $\infty$.
@@ 1.5. Goal and positioning of the paper
The goal of this paper is to provide a developed mathematical study of a numerical scheme that was proposed in our previous contribution [11]. It enters the broad family of the two-point flux approximation (TPFA) finite volume schemes [26], which has been applied to several cross-diffusion problems in recent years (see e.g. [10, 15, 36, 30, 13, 37, 9]). Our scheme mixes ideas from the so-called square-root approximation (SQRA) finite volume scheme [12] (see also [38, 33]) with some features of the Scharfetter–Gummel scheme [39, 18], especially the use of the Bernoulli function (28) in the definition of the fluxes. Note that a purely SQRA approach was also proposed in [11], but since it is overpassed by the mixed approach both in terms of accuracy and of robustness in the small Debye length (or quasi-neutral) limit, we do not develop the analysis here. Let us however stress that our whole study extends to this full-SQRA scheme at the price of very minor modifications sketched in [11].
@@
As highlighted in [11], our scheme fulfils key properties. First, it is second order accurate in space, in opposition to the scheme proposed in [7] in which upwinding leads to a mere first order consistency. Even though our scheme does not use the so-called entropy variables (the electrochemical potentials in our framework) as unknowns in opposition to the finite element scheme presented in [32], our scheme satisfies a discrete energy dissipation inequality, which should be thought of as a discrete counterpart to (9), and which is key for its numerical analysis. Keeping conservative variables as primary variables yields well-behaved non-linear solvers (we used Newton–Raphson), allowing us to consider larger time steps than for the finite element methods of [32] or for the finite volume method of [2]. Together with the stability to be used for proving the convergence of the scheme, the discrete energy dissipation inequality implies that the scheme preserves exactly the steady states of the continuous model in the sense that, at the discrete level, fluxes vanish when concentrations are of the form (19). The unique steady state of our scheme is then characterized only by the electric potential which solves a discrete counterpart of the Poisson–Boltzmann equation (20), which can be computed directly in an efficient way.
@@
In the present contribution, we first rigorously establish the convergence of the scheme when the discretization parameters (mesh size and time step) tend to 0. As in the continuous setting [31], the proof is based on compactness arguments. It strongly relies on the discrete counterpart of the energy dissipation estimate (9), following the boundedness-by-entropy method [35], and more precisely a discrete version of it [16, 37]. Uniqueness of the weak solutions is moreover established in [31] under moderate regularity assumptions thanks to the uniqueness technique of Gajewski [27, 28].
@@
We also investigate the long-time behaviour of the numerical scheme. We show in particular that the discrete solution to our scheme for the evolution problem converges as time goes to infinity towards the unique solution to the discrete steady problem. As the proof relies on LaSalle’s invariance principle, no quantitative convergence rate is derived, similarly to what is proposed in [19] in the continuous setting but in absence of electrical interactions. By establishing this result, we rigorously establish in the discrete setting the long-time behavior conjectured in [6] and recalled in Section 1.4. Numerical simulations suggest that the convergence rate strongly depends on the Debye length, but also on the initial profile $U^{0}$ (in opposition to the classical Poisson–Nernst–Planck problem without volume filling [3]).